OpenStudy (anonymous):

solve:a man left 1/2 of his money to his wife,1/6 to his daughter and remainder of 15,000 to his son. How much money did he leave?

OpenStudy (anonymous):

45000

OpenStudy (anonymous):

formula?

OpenStudy (anonymous):

1 - 1/2 - 1/6 = 2/6 2/6 -> 15000 6/6 -> 45000

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

:)

OpenStudy (ganpat):

let total money be x Therefore x/2 + x/6 + 15,000 = x Therefore x = 45,000 ...

OpenStudy (anonymous):

assume he had 45 ,000 then 1/2 of 45,000= 22,500 then 1/6 of 45000=7,500 add the totals 22,500+7,500=30,000 then add the total of wife and daughter together with that of the son=45,000